A circular pool is 34 feet in diameter and has a flag stone walk around it that is 3 feet wide. Find the area of the walk. Express the answer in terms of pie.
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

A circular pool is 34 feet in diameter and has a flag stone walk around it that is 3 feet wide. Find the area of the walk. Express the answer in terms of pie.
-------------------
Draw the picture: a circle inside a circle..
----------
Area of inner circle: pi(17)^2 sq. ft.
-----------------
Area of outer circle: pi(20)^2 sq. ft.
----------------
Area of walk = outer area - inner area = [20^2-17^2]pi sq. ft.
= 111*pi sq. ft.
======================
Cheers,
Stan H.

Answer by nerdybill(7384)

A circular pool is 34 feet in diameter and has a flag stone walk around it that is 3 feet wide. Find the area of the walk. Express the answer in terms of pie.
.
Draw a picture of the situation, this will help you "see" the problem. Imagine a donut, the "hole" is the pool and the donut is the walkway.
.
"area of circle formed by outer edge of walkway" - "area of pool" = "area of walkway"
.
"area of circle formed by outer edge of walkway":
radius = (34/2)+3 =17+3 = 20 feet
area = (pi)r^2
area = (pi)20^2
.
"area of pool":
area = (pi)r^2
area = (pi)17^2
.
"area of circle formed by outer edge of walkway" - "area of pool" = "area of walkway"
(pi)20^2 - (pi)17^2 = "area of walkway"
400(pi) - 289(pi) = "area of walkway"
111(pi) square feet = "area of walkway"